# First moments of ${\rm GL}(3)\times{\rm GL}(2)$ and ${\rm GL}(2)$ $L$-functions and their applications to subconvexity

### v1

@@ 4.1.2. Bounding $\widehat{\mathcal{T}}$
$\sum_{k\equiv 0\bmod 2}W{\left(\frac{k-1}{K}\right)}\sum_{f\in\mathcal{B}^{\ast}_{k}}\omega^{-1}_{f}\lambda_{f}(\ell)L(1/2+it,F\otimes f)$
$\hskip 65.44142pt\ll\mathcal{Y}^{\varepsilon}\left(\frac{K^{1+\varepsilon}}{\sqrt{\ell}}+\frac{\mathcal{Y}^{3/4}\ell^{1/4}}{K^{5/2-\varepsilon}}+\frac{\mathcal{Y}\sqrt{\ell}}{K^{4-\varepsilon}}+\mathcal{Y}^{9/16+\varepsilon}\ell^{7/16}+\frac{|t|^{7/4}\mathcal{Y}^{3/8+\varepsilon}{\ell}^{5/8}}{K^{5/4}}\right).$ (4.75)
By inserting the value of $\mathcal{Y}$ and a straightforward simplification, this immediately gives the estimate (1.4), as desired. This assertion of Theorem 1.4 thus follows.
4.2. Proof of Corollary 1.5
@@
At the end of this paper, it remains to complete the proof of Corollary 1.5. Indeed, after divided by a factor $\sqrt{\ell}$ on both sides of (4.2), we sum over $K^{1-\delta}\ll\ell\ll K^{1+\varepsilon}$ for any $\delta,\varepsilon>0$, obtaining
@@
$\sum_{k\equiv 0\bmod 2}W{\left(\frac{k-1}{K}\right)}\sum_{f\in\mathcal{B}^{\ast}_{k}}\omega^{-1}_{f}L(1/2,f)L(1/2+it,F\otimes g)$
@@
$\hskip 73.97733pt\ll\mathcal{Y}^{\varepsilon}\left(K+\mathcal{Y}^{9/16+\varepsilon}K^{15/16}+\frac{\mathcal{Y}^{3/4}}{K^{7/4}}+\frac{\mathcal{Y}}{K^{3-\varepsilon}}+\frac{|t|^{7/4}\mathcal{Y}^{3/8+\varepsilon}}{K^{1/8}}\right).$
This gives (1.6). The range $K^{10/9+\varepsilon}\leq t\leq K^{2-\varepsilon}$ stems from (4.74), upon taking $\ell$ such that $K^{1-\delta}\ll\ell\ll K^{1+\varepsilon}$.

### v4

@@ 3.1.2. Oscillatory case
$\sum_{k\equiv 0\bmod 2}W{\left(\frac{k-1}{K}\right)}\sum_{f\in\mathcal{B}^{\ast}_{k}}\omega^{-1}_{f}\lambda_{f}(\ell)L(1/2,F\otimes f)$
$\hskip 65.44142pt\ll\mathcal{Y}^{\varepsilon}\left(\frac{K^{1+\varepsilon}}{\sqrt{\ell}}+\frac{\mathcal{Y}^{3/4}\ell^{1/4}}{K^{5/2-\varepsilon}}+\frac{\mathcal{Y}\sqrt{\ell}}{K^{4-\varepsilon}}+\ell^{1/2+\varepsilon}+\mathcal{Y}^{1/6+\varepsilon}\right).$ (4.13)
@@
Now, after divided by a factor $\sqrt{\ell}$ on both sides of (4.1), we sum over $K^{1-\delta}\ll\ell\ll K^{1+\varepsilon}$ for any $\delta,\varepsilon>0$, obtaining
@@
$\sum_{k\equiv 0\bmod 2}W{\left(\frac{k-1}{K}\right)}\sum_{f\in\mathcal{B}^{\ast}_{k}}\omega^{-1}_{f}L(1/2,f)L(1/2,F\otimes f)\ll K^{1+\varepsilon}.$
This gives (1.6) immediately.
4.2. Proof of Theorem 1.5
At the end of this paper, it remains to complete the proof of Theorem 1.5. This can be realized by a similar argument in §4.1. Actually, by repeating the procedure in evaluating the multiple sum (4.1) with $\ell=1$ as in §4.1, one sees that the sum in (1.7) is boiled down to evaluating $\widehat{\mathcal{T}}(K,1)$, up to the error terms $\text{Err}^{\flat}(K,1)$ and $\text{Err}^{\natural}(K,1)$. We are thus allowed to deduce
$\sum_{k\equiv 0\bmod 2}W{\left(\frac{k-1}{K}\right)}\sum_{f\in\mathcal{B}^{\ast}_{k}}\omega^{-1}_{f}L(1/2,F\otimes f)=\frac{L(1,F)K}{4}\widehat{W}(0)$ $+\widehat{\mathcal{T}}(K,1)+\text{Err}^{\flat}(K,1)$
$+\text{Err}^{\natural}(K,1)+O(K^{-1/2+\varepsilon}).$
Observe that (4.12) is estimated as $O(\mathcal{Y}^{-100})$ whenever $\ell=1$, which, however, implies that $\widehat{\mathcal{T}}(K,1)$ is negligibly small in term of the magnitude. One concludes that
@@
$\sum_{k\equiv 0\bmod 2}W{\left(\frac{k-1}{K}\right)}\sum_{f\in\mathcal{B}^{\ast}_{k}}\omega^{-1}_{f}L(1/2,F\otimes f)$ $=\frac{L(1,F)K}{4}\widehat{W}(0)+O(K^{-1/4+\varepsilon}),$
as desired.